# Time Distributed Deep Learning models for Purely Exogenous Forecasting. Application to Water Table Depth Prediction using Weather Image Time Series

### v1

@@ 1.2 Case Study Description
Groundwater resources in Italy are even more exploited than European average statistics. In fact, in Italy 85% of drinking water [68] is extracted from groundwater sources. In Piemonte, an administrative Region in the north-west of Italy, nearly half of total water abstractions are for the agriculture sector [69, 70] which extensively adopts irrigation to meet the water needs of crops. It is estimated that 83% of irrigable lands are effectively irrigated [71], a fact that states the limited use of seasonal precipitation as a direct source of water in place of human abstractions.
Piemonte is a very heterogeneous region from the geographical point of view, with Alps near the western, northern and southern borders, hills extending from the centre to south-est, and plains which cover an area from the Cuneo Province in the south-west to the upper-central part in Vercelli and Novara Provinces. In this context is very difficult to analyse groundwater resources, also because of the intensive agricultural activities. In fact, [72] analyzed aquifer recharge in the Piemonte Alpine zone in the north-west, and authors highlighted the difficulties of finding a general trend among different areas, and the results of the analysis are very context-dependent. For this reason, we decided to focus on a specific catchment named Grana-Maira, located in the Cuneo Province in the south-west of Piemonte (Figure 1). The Grana-Maira catchment takes its name from the two rivers (Grana and Maira) which originate in the Alps in the western part of the catchment. The elevation of the catchment terrain decreases from west to northeast, i.e. from the mountain to the plain.
Figure 1: Piemonte Region with administrative Province (bold names). Zoom on the Grana-Maira selected catchment (in magenta) with the three water table sensors. The digital elevation model (DTM) is represented in meters [m] through a grey logarithmic scale.
@@ 1.2.1 Data
We retrieved three time series of the water table depth from sensors in the municipalities of Vottignasco, Savigliano, and Racconigi [^3] More sensors were available in the catchment, however other sensors measured shorter series or had longer missing periods. For this reason we decided to get the data by only the mentioned three sensors.. These sensors are part of the measurement network of the Regional Environmental Agency (ARPA [^4] In Italian Agenzia Regionale per la Protezione Ambientale.) and are freely available by request. Table 1 reports different information and summary statistics for the three series, and Figure 2 shows the weekly average series for the three sensors. This figure highlights a problem of missing data and irregularities between the series. In fact, some huge missing periods are present and, furthermore, these gaps are not the same for the three series. Given the large extension of some of these missing periods, it was considered more sensible not to impute any data, but to try to build deep learning models able to learn without making imputation.
For the present work, weather raster data of total precipitation, maximum, and minimum temperature at 0.125° spatial resolution were retrieved from [73] and solely used as the input of the proposed models. We decided to not include an autoregressive component (i.e. past values of water table depth as an additional input) because groundwater data are released on a semester basis, then using an autoregressive component would make the proposed models unusable at present time because of the lack of the recent water table data. Instead, weather data are updated daily without any missing values, then for each time step $t$ three images at 0.125° resolution are available, one for each variable. Thus, for each water table depth point to be predicted, it is possible to construct a video made of frame with three channels, namely total precipitation, maximum and minimum temperature.
Figure 2: a) Water Table Depth weekly time series measured by sensor Vottignasco 00425010001 b) Water Table Depth weekly time series measured by sensor Savigliano 00421510001 c) Water Table Depth weekly time series measured by sensor Racconigi 00417910001. All the plots have the y-axis reversed (lower value on top) because in this way it should be easier to interpret the water table depth: if it increases it means a decrease in water resources stored in the ground.
Table 1: Water Table Depth time series statistics. Sensor column shows the municipality and the identification codes of the sensors; $\mu$ represents the global mean; $\sigma$ represent the global standard deviation; Observation represents the total number of observations; First and Last data represent the dates of the first and last measurement respectively.
@@ 2 Methods
This work has aimed to train local models for each sensor independently of the others. This is because it is consistent with the literature (see for example [25, 27, 21]) but also because of the irregularities between the series explained in Section 1.2. In fact, if one wanted to train a global model using the data from all the sensors, it would require getting only the data from non-missing dates available jointly to all three sensors, discarding in this way a lot of information. Furthermore, as shown in Section 1.2, the three series show different dynamic behaviour and a general global model could be too restrictive and of scarce utility for the domain application.
The local models we have developed take as input a multivariate and multidimensional time series, i.e. a video, of time length $T$, spatial extent $H\times W$ (i.e. height and width of each frame), and $P$ weather features. Each local model forecasts the water table depth at a weekly time step $t$ in a sliding windows fashion. Formally, a local model has to learn the relation $f$ which links $y_{t}=f(X_{t-1}^{H,W,P,T})+\epsilon_{t}$, where $\epsilon$ is the irreducible error term, and the input $X_{\bullet}^{H,W,P,T}$ is defined as the multivariate image sequence $X_{t}^{H,W,P,T}=\{X(h;w;p;\tau):h\in[1;H],w\in[1;W],p\in[1;P],\tau\in[t-T+1;t]\}$, where $X$ denotes the tensor composed by the full length image time series. Given that groundwater phenomena $y$ could have a very long memory, we have used a very long input weather image time series length, setting T to 104 (i.e. 2 years in weekly terms) and letting the DL models use the most relevant past information.
As already stated, the proposed models are made of two modules. The first module is a Time Distributed CNN (TDC) whose structure is identical in the two models. The TDC is responsible for learning spatial relations and it outputs a Time Distributed Hidden Representation of the input image time series (Figure 3), i.e. it encodes each frame of the video into a vector of dimension $D$; in other words, it extracts a classical multivariate (of $D$ variables) time series from the input. The second module, the Sequential Module, is what characterizes the two different models called TDC-LSTM and TDC-UnPWaveNet. In the following, the building boxes of the two models are explained, especially the modification and novelties carried out in the development of the UnPWaveNet, as the Channel Distributed layer.
Figure 3: Modeling Pipeline.
@@ 2.1.1 Time Distributed Layers
Generally speaking, inside a standard Neural Network, a hidden layer is made of many neurons, each of which is responsible of computing an affine transformation of the input and applying a non-linear activation function $g$. Formally $A_{k,l}=g(W_{k,l}^{\intercal}A_{l-1}+\beta_{k,l})$ where $A_{k,l}$ is the output of a general neuron $k$ in layer $l$, and $A_{l-1}$ is the output matrix of the layer $l-1$; $W$ and $\beta$ contain the parameters to be learned. It is possible to substitute the ”simple” neuron with more complex operations, still made of neurons, but encapsulated in a so-called cell, while the formal neurons inside the cell are referred to as ”units”. An exemplification concerning sequential data could be RNNs, and in particular, the LSTM layer, which is made of as many cells as the number of elements in the input sequence (or time step in the case of temporal data). Every cell is responsible for extracting long and short-term temporal dependencies and passing this information to the subsequent cell (through recurrent connection) [35, 74]. However, in the LSTM layer, and in RNN in general, the weights used by neurons in a cell are the same for every cell in the layer. In other words, each element of the input sequence is processed by a cell with the same parameters, what differs between the cells of an LSTM layer are the inputs of every cell.
@@
The concept of applying the same set of operations to every element of the input sequence is not applied only for RNNs, but it is a general way of proceeding also in other architectures. A layer that works in this way is usually referred to as a Time Distributed (TD) layer. Not by chance, in Keras [^5] Open Source Python library for developing neural network models. there is a specific layer-class named TimeDistributed which applies the same set of operations to each element of the input sequence. Figure 4(a) represents a general TD layer made of a simple fully connected cell applied on a multivariate time series $Z$ of length $T$ with $C$ variables. In that case, the TD layer acts time-step by time-step transforming the vector with $C$ elements into a vector of $C^{*}$ elements, in which $C^{*}$ is defined by the number of neurons of the fully connected cell. If $C^{*}>C$ the TD layer will dilate the channel dimension of each time step, reversely if $C^{*}<C$ (as in Figure 4(a)) the TD layer will squeeze the channel dimension. It is relevant to point out that with a TD layer, the ordering and the length of the input sequence $Z$ are untouched and maintained also in the output sequence $Z^{\prime}$. It is possible to develop TD layers with other types of cells than fully connected, and, as we already discussed, one of the most used TD networks for multidimensional sequences (e.g. video) is the TD CNN [54, 55, 75]. In the case of TD CNN, the same CNN is applied on each frame of the video extracting spatial features. In Section 2.3.1 we describe the implementation of our TDC module based on TD CNN.
@@ 2.1.2 Channel Distributed Layers
While analyzing the TD layers, a question caught our attention: Why not apply the behaviour of TD layer to channels (i.e. variables) instead of time? This implies processing each channel individually and performing computations on the temporal dimension. In this fashion, the channel-wise information is preserved, while the sequential (temporal) information is processed with the same cell for each channel. This brought us to develop the Channel Distributed (CD) cell-based layers that, as explained in the following, have been adopted in the development of the UnPWaveNet architecure.
Instead of looking at the multivariate time series $Z$ as a series of $T$ time-indexed vectors each with $C$ element, it is possible to interpret it as a set of $C$ univariate time series. Then, taking a univariate series $c_{j}$ with $j\in[1;C]$, it is feasible to feed the $c_{j}$ series into a cell and apply the same cell $\forall c_{j},j\in[1;C]$. Figure 4(b) represents a CD layer with a fully connected cell. If the number of neurons $T^{*}$ in the cell is less than the number of elements in the input univariate series (i.e. $T<T^{*}$), the CD layer compresses the time dimension of the input sequence, leaving untouched the channel dimension which still contains $C$ variables. Reversely, if $T>T*$ the CD layer will expand the time dimension. This means that the CD layer squeezes or dilates the time dimension of all channels with the same cell; that is exactly what the TD layer does but acting instead on the time dimension. As for the TD layer, the cell could take any form. However, in our application we found the fully connected cell to produce already satisfying results in the CD layer of the UnPWaveNet.
@@ 2.2 WaveNet & UnPWaveNet
In the last year, many studies have tried to develop new convolutional models for temporal-sequential data tasks to compete with RNN[37, 38, 39, 42, 43, 76]. Most of these works are based on the dilated convolution, which enables the exponential expansion of the receptive field of the network over the input sequence (i.e. look far away in the past of the input series) [45]. WaveNet [46] is exactly based on this concept, and it also integrates a causal constraint using causal padding. The causal padding makes every element of the output sequence to depend only on current and past input data, and not on the future. Furthermore, it forces the output to have the same length as the input sequence. All this is achieved by sliding the convolution operations from right (more recent values) to left (older values) and adding zeros to the left of the input [^6] For more details on the dilated convolution and causal passing look at [45, 37, 46, 39, 40]. In [46], this processing is named dilated causal convolution, shown in Figure 5(a). Even if the WaveNet model was designed for audio generation, thanks to its ability to handle temporal data, it has been applied with remarkable results also to other tasks, among which financial data [40] and hydrology [77]. Each layer of the WaveNet (Figure 5(b)) consists of a dilated causal convolution, a gated activation unit [78] [^7] Gate activation units are also employed in the LSTM cell, in which are named gates. Other works employed and improved this type of activation [79] obtaining better results than using the classic ReLU., and a 1x1 convolution [^8] The 1x1 convolution is convolution with a kernel of dimension 1 which acts as a bottleneck squeezing the channel dimension [31]. This is equivalent to a TD fully connected layer with as much neurons as the number of filter of the convolution.. What makes WaveNet very flexible are the residual connections over the dilated causal convolutions, and the skip connections which concatenate the result of every 1x1 convolution letting the gradient flow easily over the network.
@@
Even if the WaveNet and other convolutional-based networks have brought astonishing results, they have been usually applied for many-to-many tasks, and in general to predict sequences of the same length of the input. We have restructured the WaveNet architecture to predict output sequences completely shifted in the future and shorter than the input sequences. In this case, the causal constraint is no longer needed, because the output is completely in the future and, thus each output element should depend on every input element. Consequently, in such a case, it is possible to drop the causal padding implemented in the WaveNet and let the temporal dimension be squeezed layer after layer. Figure 6(a) represents dilated convolution without the causal padding, here named as unpadded dilated convolution, whose receptive field $r_{l}$ for a layer $l$ with a filter of dimension $K$ is defined by equation 1. It is worth noting that without the causal constraint, the first element of the output sequence of a layer $l$ has a receptive field $r_{l}$ which covers the first $r_{l}$ elements of the input sequence, while the last element of the output sequence has a receptive field which covers the last $r_{l}$ elements of the input sequence. When instead the causal constraint is enforced every element of the output sequence of a layer $l$ has a receptive field which covers up to $r_{l}$ first (i.e. past) element of the input sequence.
@@
Two problems arise when removing the causal padding constraint from the initial network architecture: it is a) no longer possible to add residual connections; and b) no longer possible to concatenate skip connections. The cause for both problems is the different dimensions of the input and output sequences of each layer, which without the padding are no longer maintained over the network. To solve problem a), we apply a pooling average layer to meet the dimension of the 1x1 convolution. While to solve problem b) we have adopted our proposed Channel Distributed layer with a fully connected cell, which transforms the output sequences from 1x1 convolutions into new sequences with the same length equal to the length of the requested output. It is relevant to point out that the CD layer with the fully connected cell does not respect any causal constraint, in fact, each output element of the employed CD layers depends on all the input elements provided by the 1x1 convolution. Figure 6 shows the architecture of our proposed version of the WaveNet, here renamed UnPWaveNet (unpadded WaveNet). With respect to the original implementation (Figure 5(b)), the UnPWaveNet use only one 1x1 convolution layer instead of two after the skip concatenation. This is because the CD layers already apply some transformation to the skip connections, and thus it was sufficient to achieve satisfactory results saving parameters.

### v2

@@ 1.2 Case Study Description
Groundwater resources in Italy are even more exploited than the European average. Indeed, 85% of drinking water in Italy comes from groundwater sources [74]. In Piedmont, an administrative region in the north-west of Italy, almost half of the total water abstracted is used by the agricultural sector, which makes extensive use of irrigation to meet the water needs of crops [75, 76]. It is estimated that 83% of irrigable land is effectively irrigated, a fact that indicates the limited use of seasonal rainfall as a direct source of water instead of human abstractions [77].
Piedmont is a very heterogeneous region from the geographical point of view, with Alps near the western, northern, and southern borders, hills extending from the center to south-est, and plains that cover an area from the Cuneo Province in the south-west to the upper-central part in Vercelli and Novara Provinces. In this context is very difficult to analyze groundwater resources, also because of the intensive agricultural activities. Indeed, the authors in [78] analyzed aquifer recharge in the Piedmont Alpine zone in the north-west and they highlighted the difficulty of finding a general trend between different areas, in other words, the results of their analysis are very context-dependent. For this reason, we decided to focus on a specific catchment called Grana-Maira, located in the Cuneo Province in the southwest of Piedmont (Figure 1). The Grana-Maira basin takes its name from the two rivers (Grana and Maira) that originate in the Alps in the western part of the basin. The altitude of the catchment area decreases from west to northeast, i.e. from mountain to plain.
Figure 1: Piedmont Region with administrative Province (bold names). Zoom on the Grana-Maira selected catchment (in magenta) with the three water table sensors. The digital elevation model (DTM) is represented in meters [m] through a grey logarithmic scale.
@@ 1.2.1 Data
We retrieved three time series of the water table depth from sensors in the municipalities of Vottignasco, Savigliano, and Racconigi. These sensors are part of the measurement network of the Regional Environmental Agency (ARPA [^2] In Italian Agenzia Regionale per la Protezione Ambientale.) and are freely available by request. Table 1 reports different information and summary statistics for the three series, and Figure 2 shows the weekly average series for the three sensors. This figure highlights a problem of missing data and irregularities between the series. Indeed, some huge missing periods are present and, furthermore, these gaps are not the same for the three series. Given the large extension of some of these missing periods, it was considered more sensible not to perform imputation, but to let the deep learning models learn from the available data.
For the present work, weather raster data of total precipitation, maximum, and minimum temperature at 0.125° spatial resolution were retrieved from [79] and solely used as the input of the proposed models. We decided to not include an autoregressive component (i.e. past values of water table depth as an additional input) because groundwater data are released on a semester basis, then using an autoregressive component would make the proposed models unusable at present time because of the lack of the recent water table data. Instead, weather data are updated daily without any missing values. Thus, for each water table depth point to be predicted, it is possible to construct a video made of frames each one with three channels, namely total precipitation, and maximum and minimum temperature.
Figure 2: Water Table Depth weekly time series measured by sensors a) Vottignasco 00425010001, b) Savigliano 00421510001, and c) Racconigi 00417910001. All the plots have the y-axis reversed (lower value on top) because in this way it should be easier to interpret the water table depth: if it increases it means a decrease in water resources stored in the ground.
Table 1: Water Table Depth time series statistics. The sensor column shows the municipality and the identification codes of the sensors; $\mu$ represents the global mean; $\sigma$ represents the global standard deviation; Observation represents the total number of observations; First and Last data represent the dates of the first and last measurement respectively.
@@ 2 Methods
This work has aimed to train local models for each sensor independently of the others. This is because it is consistent with the literature (see for example [31, 33, 22]) but also because of the irregularities between the series explained in Section 1.2. Indeed, if one wanted to train a global model using the data from all the sensors, it would require getting only the data from non-missing dates available jointly to all three sensors, discarding in this way a lot of information. Furthermore, as shown in Section 1.2, the three series show different dynamic behavior and a general global model could be too restrictive and of scarce utility for the domain application.
Our proposed modeling pipeline is described in Figure 3, and it consists in black box local models that take as input a multivariate and multidimensional time series, i.e. a video $X_{t}^{H,W,P,T}=\{x(h;w;p;\tau)\in\mathbb{R}\,|\,h\in[1;H],w\in[1;W],p\in[1;P],\tau\in[t-T+1;t]\}$, of spatial extent $H\times W$ (i.e. height and width of each frame), time length $T$, and $P$ weather features. Each local model forecasts the water table depth at a weekly time step $t$ in a sliding window fashion. Formally, a local model has to learn the relation $f$ which links $y_{t}=f(X_{t-1}^{H,W,P,T})+\epsilon_{t}$, where $\epsilon$ is the irreducible error term. Given that groundwater phenomena $y$ could have a very long memory, we have used a very long input weather image time series length, setting T to 104 (i.e. 2 years in weekly terms) and letting the DL models use the most relevant past information.
The general structure of the proposed black box models consists of two modules (Figure 3). The first module is a Time Distributed CNN (TDC) whose structure is identical to the two models. The TDC is responsible for learning spatial relations and it outputs a Time Distributed Hidden Representation of the input image time series, i.e. it encodes each frame of the video into a vector of dimension $D$; in other words, it extracts a classical multivariate (of $D$ variables) time series from the input. The second module, the Sequential Module, is what characterizes the two different models called TDC-LSTM and TDC-UnPWaveNet. In the following, the building boxes of the two models are explained, especially the modification and novelties carried out in the development of the UnPWaveNet, as the Channel Distributed layer.
Figure 3: Proposed modeling pipeline with the general structure of the black box local models.
@@ 2.1.1 Time Distributed Layers
Generally speaking, inside a standard Neural Network, a hidden layer is made of many neurons, each of which is responsible for computing an affine transformation of the input and applying a non-linear activation function $g$. Formally $A_{k,l}=g(W_{k,l}^{\intercal}A_{l-1}+\beta_{k,l})$ where $A_{k,l}$ is the output of a general neuron $k$ in layer $l$, and $A_{l-1}$ is the output matrix of the layer $l-1$; $W$ and $\beta$ contain the parameters to be learned. It is possible to substitute the ”simple” neuron with more complex operations, still made of neurons, but encapsulated in a so-called cell, while the formal neurons inside the cell are referred to as ”units”. An exemplification concerning sequential data could be RNNs, and in particular, the LSTM layer, which is made of as many cells as the number of elements in the input sequence (or time step in the case of temporal data). Every cell is responsible for extracting long and short-term temporal dependencies and passing this information to the subsequent cell (through recurrent connection) [41, 80]. However, in the LSTM layer, and in RNN in general, the weights used by neurons in a cell are the same for every cell in the layer. In other words, each element of the input sequence is processed by a cell with the same parameters, what differs between the cells of an LSTM layer are the inputs of every cell.
@@
The concept of applying the same set of operations to every element of the input sequence is not applied only in RNNs, but it is a general way of proceeding also in other architectures. A layer that works in this way is usually referred to as a Time Distributed (TD) layer. Not by chance, in Keras [^3] Open Source Python library for developing neural network models https://keras.io/. there is a specific layer-class named TimeDistributed which applies the same set of operations to each element of the input sequence. Figure 4(a) represents a general TD layer made of a simple fully connected cell applied on a multivariate time series $Z$ of length $T$ with $C$ variables, i.e. $Z=\{Z_{\tau}\in\mathbb{R}^{C}|\ \tau\in[t-T,t-1]$}. In Figure 4(a) the TD layer acts time-step by time-step transforming each vector $Z_{\tau}\in\mathbb{R}^{C}$ into a vector $Z^{\prime}_{\tau}\in\mathbb{R}^{C^{*}}$, in which $C^{*}$ is defined by the number of neurons of the fully connected cell. If $C^{*}>C$ the TD layer will dilate the channel dimension of each time step, reversely if $C^{*}<C$ (as in Figure 4(a)) the TD layer will squeeze the channel dimension. It is relevant to point out that with a TD layer, the ordering and the length of the input sequence $Z$ are untouched and maintained also in the output sequence $Z^{\prime}$. It is possible to develop TD layers with other types of cells than fully connected, and, as we already discussed, one of the most used TD networks for multidimensional sequences (e.g. video) is the TD CNN [60, 61, 81]. In the case of TD CNN, the same CNN is applied on each frame of the video extracting spatial features. In Section 2.3.1 we describe the implementation of our TDC module based on TD CNN.
@@ 2.1.2 Channel Distributed Layers
While analyzing the TD layers, a question caught our attention: Why not apply the behavior of TD layer to channels (i.e. variables) instead of time? This implies processing each channel individually and performing computations on the temporal dimension. In this fashion, the channel-wise information is preserved, while the sequential (temporal) information is processed with the same cell for each channel. This brought us to develop the Channel Distributed (CD) cell-based layers that, as explained in the following, have been adopted in the development of the UnPWaveNet architecture.
Instead of looking at the multivariate time series $Z$ as a series of $T$ time-indexed vectors each with $C$ element, it is possible to interpret it as a set of $C$ univariate time series $Z=\{c_{j}\in\mathbb{R}^{T}|\ j\in[1;C]$}. Figure 4(b) shows how a CD layer with a fully connected cell works. More specifically, in Figure 4(b) the CD layer acts channel by channel, transforming each univariate time series $c_{j}\in\mathbb{R}^{T}$ into a new series $c_{j}^{\prime}\in\mathbb{R}^{T^{*}}$ $\forall\ j\in[1;C]$. If the number of neurons $T^{*}$ in the cell is less than the number of elements in the input univariate series (i.e. $T<T^{*}$), the CD layer compresses the time dimension of the input sequence, leaving untouched the channel dimension which still contains $C$ variables. Reversely, if $T>T*$ the CD layer will expand the time dimension. This means that the CD layer squeezes or dilates the time dimension of all channels with the same cell; that is exactly what the TD layer does but acting instead on the time dimension. As for the TD layer, the cell could take any form. However, in our application we found the fully connected cell to produce already satisfying results in the CD layer of the UnPWaveNet.
@@ 2.2 WaveNet & UnPWaveNet
In the last year, many studies have tried to develop new convolutional models for temporal-sequential data tasks to compete with RNN[43, 44, 45, 48, 49, 82]. Most of these works are based on dilated convolution, which enables the exponential expansion of the receptive field of the network over the input sequence (i.e. look far away in the past of the input series) [51]. WaveNet [52] is exactly based on this concept, and it also integrates a causal constraint using causal padding. The causal padding makes every element of the output sequence to depend only on current and past input data, and not on the future. Furthermore, it forces the output to have the same length as the input sequence. All this is achieved by sliding the convolution operations from right (more recent values) to left (older values) and adding zeros to the left of the input [^4] For more details on the dilated convolution and causal passing look at [51, 43, 52, 45, 46]. In [52], this processing is named dilated causal convolution, shown in Figure 5(a). Even if the WaveNet model was designed for audio generation, thanks to its ability to handle temporal data, it has been applied with remarkable results also to other tasks, among which financial data [46] and hydrology [83]. Each layer of the WaveNet (Figure 5(b)) consists of a dilated causal convolution, a gated activation unit [84] [^5] Gate activation units are also employed in the LSTM cell, in which are named gates. Other works employed and improved this type of activation [85] obtaining better results than using the classic ReLU., and a 1x1 convolution [^6] The 1x1 convolution is convolution with a kernel of dimension 1 which acts as a bottleneck squeezing the channel dimension [37]. This is equivalent to a TD fully connected layer with as many neurons as the number of filters of the convolution.. What makes WaveNet very flexible are the residual connections over the dilated causal convolutions and the skip connections which concatenate the result of every 1x1 convolution letting the gradient flow easily over the network.
@@
Even if the WaveNet and other convolutional-based networks have brought astonishing results, they have been usually applied for many-to-many tasks, and in general to predict sequences of the same length of the input. We have restructured the WaveNet architecture to predict output sequences completely shifted in the future and shorter than the input sequences. In this case, the causal constraint is no longer needed, because the output is completely in the future and, thus each output element should depend on every input element. Consequently, in such a case, it is possible to drop the causal padding implemented in the WaveNet and let the temporal dimension be squeezed layer after layer. Figure 6(a) represents dilated convolution without the causal padding, here named as unpadded dilated convolution, whose receptive field $r_{l}$ for a layer $l$ with a filter of dimension $K$ is defined by Equation 1. It is worth noting that without the causal constraint, the first element of the output sequence of a layer $l$ has a receptive field $r_{l}$ which covers the first $r_{l}$ elements of the input sequence, while the last element of the output sequence has a receptive field which covers the last $r_{l}$ elements of the input sequence. Instead, when the causal constraint is enforced, every element of the output sequence of a layer $l$ has a receptive field that covers up to $r_{l}$ first (i.e. past) element of the input sequence.
@@
Two problems arise when removing the causal padding constraint from the initial network architecture: it is a) no longer possible to add residual connections; and b) no longer possible to concatenate skip connections. The cause for both problems is the different dimensions of the input and output sequences of each layer, which without the padding are no longer maintained over the network. To solve the problem a), we apply a pooling average layer to meet the dimension of the 1x1 convolution. While to solve problem b) we have adopted our proposed Channel Distributed layer with a fully connected cell, which transforms the output sequences from 1x1 convolutions into new sequences with the same length equal to the length of the requested output. Figure 6(b) shows the architecture of our proposed version of the WaveNet, here renamed UnPWaveNet (unpadded WaveNet). With respect to the original implementation (Figure 5(b)), the UnPWaveNet uses only one 1x1 convolution layer instead of two after the skip concatenation. This is because the CD layers already apply some transformation to the skip connections, and thus it was sufficient to achieve satisfactory results saving parameters.